OpenStudy (anonymous):

Sorry I asked a question wrong earlier. =] A room has an area of 189 square ft. If the room is square how far would an ant walk to get AROUND the perimeter?

OpenStudy (diyadiya):

54.99

OpenStudy (saifoo.khan):

Area = b^2 189 = b^2 \[b=\sqrt{189}\] b = 13.74ft this means that one side of the room is, 13.74

OpenStudy (saifoo.khan):

Now, multiply this by 4, 13.74 * 4 = 54.95ft.

OpenStudy (anonymous):

Well, if a is the side of the square then,\[a^2=189\]\[a=13.8\]So the perimeter is equal to\[P=4a=4\times13.8=55.2\]